OpenStudy (anonymous):

Find the solution of the following differential equation that passes through (-1,-1). You may use separation of variables. dy/dt = [ 2t(y+1) ] / [y] This is what I have after separation of variables and integration of both sides: y - ln(y+1) = t^2 +c Plugging in (-1,-1) would not work as ln(0) does not exist. Is there another way to solve this problem? Thanks in advance!

myininaya (myininaya):

this is weird my solution also has ln(y+1)

myininaya (myininaya):

it seems like they would give points for which the function is defined

OpenStudy (anonymous):

my solution doesn't.

myininaya (myininaya):

what did you get?

myininaya (myininaya):

it has to be that it is undefined. i feel confident in my answer

myininaya (myininaya):

let me ask loiskin though one sec

myininaya (myininaya):

oh he left nvm

OpenStudy (anonymous):

hint: it is NON-LINEAR ordinary diffeq

myininaya (myininaya):

okay we have y dy/(y+1)=2t dt. Ingratiating both sides gives us y+1-ln(y+1)=t^2/2+C

myininaya (myininaya):

for the y part, I used a substitution y+1=u so dy=du and we we can find y=u-1 so we have int([u-1]/u, u)=u-lnu= y+1-ln(y+1)

myininaya (myininaya):

therefore, the function does not exist at (-1,-1)

myininaya (myininaya):

oops the t side is both to have t^2+C, that was a type-0

myininaya (myininaya):

suppose to not both to lol stupid type-0s

OpenStudy (anonymous):

nope. you didn't finish the problem.

myininaya (myininaya):

ln(y+1) does not exist for y=-1

myininaya (myininaya):

(-1,-1) is not in the domain or range

OpenStudy (anonymous):

keep going.

myininaya (myininaya):

just tell me what you are thinking. I'm not the one who needs this problem done. lol

myininaya (myininaya):

that guy left along time ago it looks like

myininaya (myininaya):

did you want me to say there isn't a constant that exist with the initial condition y(-1)=-1. Is this what you wanted me to say?

OpenStudy (anonymous):

I got that answer, but I'm looking at something.

OpenStudy (anonymous):

Hmm, I have to say, given that this is a first order differential equation, and you've found a solution, and solutions are unique, if the solution satisfies the d.e., then that's it! Couldn't it be the case you have a typo. for the BC's?

myininaya (myininaya):

he already left. I just wanted to make sure I said what I needed to about the solution since quantish felt I was wrong about the solution or something. I'm not really sure what quantish thinks. Thanks for looking lokisan.

OpenStudy (anonymous):

Do you have maple, mathmatica, or matlab?

myininaya (myininaya):

i have maple

OpenStudy (anonymous):

I've run a few different things in Wolfram Alpha, and I'm pretty sure no solutions exist.

OpenStudy (anonymous):

so wolframalpha said no solution exist?

OpenStudy (anonymous):

Wolfram Alpha just confirmed what I worked out above. I tried to get it to repeat the process and solve given the initial condition and it reports "No solutions exist". http://www.wolframalpha.com/input/?i=-1+%3D+-W%28-e^%28-x-1-1%29%29-1+

OpenStudy (anonymous):

There's something wrong with your question, myininaya.

myininaya (myininaya):

it isn't mine lol

myininaya (myininaya):

i was just worried about it because it makes me worry when people disagree with me

OpenStudy (anonymous):

Oh, well, I think you have the weight of evidence on your side ;)

myininaya (myininaya):

ok thanks for supporting my solution.

OpenStudy (anonymous):

No problems.

OpenStudy (anonymous):

I played around with that too and kept getting similar answers. I'll have to ask my professor. :/